OpenStudy (anonymous):

Given that p = mq + r, express q in terms of p, m, and r. Given that y=3x-2, express x in terms of y.

OpenStudy (anonymous):

Express q in terms of other letters: mq = p - r -> q = (p - r)/m Express x in terms of other letters: 3x = y + 2 -> x = ...